Jackson, Inc. has been in business for many years. Retained earnings on January 1, 2017, is...

  1. Jackson, Inc. has been in business for many years. Retained earnings on January 1, 2017, is $215,800. The following information is available for the first two months of 2017:

    January February
    Revenues $83,000 $96,000
    Expenses 89,000 82,000
    Dividends paid 0 5,000

    Required:

    Prepare a statement of retained earnings for the month ended February 28, 2017.

    Jackson, Inc.
    Statement of Retained Earnings
    For the Month Ended February 28, 2017
    $
    $

Check My Work5 more C
Answer #1

Addition to Retained earnings = revenue - expenses- dividend paid

Jan = 83000-89000 =-$6,000

Feb = 96000 - 82000 =$14000

Statement of retained earnings

Beginning balance 215,800

Addition during Jan -6000

Addition during Feb 14000

Ending balance of retained earnings =$223,800